The price elasticity of demand for MedPride Powder-Free Nitrile Exam Gloves, Medium, Box/100, has been estimated to be about −2.0. Data from the MedPride company for 2019 shows that 80 (thousand) boxes per month of the gloves were sold that year. In 2019, the average retail price for a box of the gloves was $16.00. Using this information, derive the linear demand function for MedPride Powder-Free Nitrile Exam Gloves in thousands of boxes in 2019.  
Answer #1

Price Elasticity of demand do . P api Slope of demand curue = -a - o poaches Ther, dp = -1/0 which means Supe of demand curue